OpenStudy (goformit100):

Prove that 3999991 is not prime.

OpenStudy (goformit100):

@TheForbiddenFollower @Thoughts

OpenStudy (anonymous):

\(3999991=4000000−9=2000^2−3^2\)

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

Use : \(a^2 - b^2 = (a+b)(a-b) \) to get the factors of the given number. (Proved : Its not prime)

OpenStudy (goformit100):

ok I got the idea :) Thank you sir